March 25, 2015 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, March 25, 2015
11:00 am, EDT
JLab: CEBAF Center, F326

Connection Using Bluejeans

  1. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 630804895.
  2. To join via a Web Browser, go to the page [1]
  3. To join via phone, use one of the following numbers and the Conference ID: 630804895
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4
  5. More information on connecting to bluejeans is available.


  1. Announcements
  2. Subdetector Reports
  3. Detector Matching (Paul): Run Browser
  4. Spring Calibration Plans
  5. Simulations
  6. AOT


Attending: Sean (NU); Simon, Kei, Paul, Lubomir, Will M., Alex B., Justin, Beni, Mike S., Mark D., Elton (JLab); Naomi (CMU); Noemi (Regina)

Subdetector Reports

  • FCAL - No news. Sean will follow up and talk to Matt & John Z. about their skimming needs
  • BCAL - When Will looks at 900 MeV showers, he gets a pi0 resolution of 9.5 MeV, although it's statistics limited. Mark D. is looking at the corrections in more detail.
  • CDC - Naomi is making progress with her drift time fitting procedure and is working on identifying 50-100 "good straws" to monitor.
  • FDC - Lubomir is finishing up his alignment calculations. He has given constants for FDC packages 2-4 to Simon, who is making the necessary changes to the reconstruction code and CCDB. The TDC times are not being currently used in the reconstruction; Lubomir and Simon will discuss this.
    • Naomi mentioned that David has implemented her new fADC125 algorithm in the offline software, and sent out an email with some instructions on how to enable it. She is still making some final checks, and the effects of this algorithm will need to be studied.
  • TOF - Sasha is still working with his algorithm, and running into problems dealing with the low signal/background and small statistics of the bars far from the beamline. After the last calibration meeting, Sean, Mark I., Simon, and Beni met together to discuss various other possible calibration methods. Beni had determined some per-channel offsets and gave them to Simon who is studying them. Simon is also working on determining the effective velocities for each bar using hits matched to tracks.
    • Sean looked at the relative calibration of the raw ADC/TDC hit times and found that some ranges were well calibrated, but the were frequent shifts in both systems. Since these time offsets can change in an uncontrolled fashion, it is important to have a prompt calibration of them. Mike S. is running his time calibration code over last Fall's data and should have updated constants soon.
  • ST - Eric is making progress in updating the timewalk and propogation corrections. He realized there was a bug in his procedure while on shift this morning and is working to fix it.
  • TAGH - Nathan was on shift during the meeting, but continues to work on aligning the tagger and PS times.
  • TAGM - Alex is adjusting the bias voltages by looking at the single-pixel counts in order to try to recover some of the channels with low efficiency.

Detector Matching

Simon added some information to look better at some of the details of track matching, and Paul walked us through some of the new plots he had added to study matching efficiencies. These efficiencies are the combination of many different effects: detector efficiency, overall reconstruction efficiency and quality, ghost tracks, etc. Paul is still studying the details - one substantial current problem is that there are a lot of bad/ghost tracks that look like they are of good quality.

  • BCAL - The matching requires a hit in the start counter, and the rate is low, ~20-25%
    • Elton had been looking at some shower reconstruction, and one major issue he noticed is that the shower reconstruction assumes that the particle that generates it is a neutral particle, and is not as good for charged particles which bend (sometimes by large amounts)
    • It might be better to look at matching rates in some 3D projection as opposed to just the 2D surface of the BCAL.
    • It would be useful to check to make sure that tracks have hits in the outer layers of the CDC. The CDC/FDC transition region is difficult, as always
  • FCAL - The matching requires a hit in the SC or TOF. The matching is ~70%; junk tracks could be the main culprit here? The decrease in rates at large radius is not understood.
  • SC - The matching requires a hit in the BCAL, FCAL, or TOF. The matching is nearly 100% at large angle, dropping to ~70% in the nose region. Reconstruction in the nose is more difficult, but again there could be a bias due to the large number of junk tracks/beam background in the forward direction. There might be a bias at large angle due to the BCAL hit matching as well.
  • TOF - The results are complicated, and reflect mainly the fact that there are no per-channel calibrations in the CCDB at the moment.

Spring Calibration Plans

We discussed the proposed plans below, and Sean made several updates to it. Elton strongly suggested that we should try to calculate a normalization for our data set this fall, as it will force us to understand various features of the experiment. There was a long discussion about how we might accomplish this. One possibility would be to use the PS, which would require developments in its reconstruction and simulation. One could also use a well-known physics process, like rho photoproduction, which requires extensive understanding of our acceptance and efficiencies. These options will be studied, but it was generally felt that an additional [clever] procedure should be developed.

Mark I. pointed out that there is a well-developed procedure used in Hall B, but it requires a working TAC and the ability to trigger on the tagger, and it is not clear if we will have either during this run.


A few things still need to be looked at before launching a new set of simulations. Sean will organize the next launch when he is at JLab in April.

Proposed Spring Plans

  1. Short-term (few days)
    1. Realign detector times
    2. Compare detector performance to fall run
  2. Medium-term (~1-2 month)
    1. Check stability of existing calibrations
    2. Improve timing calibrations
      1. Continue TDC commissioning
      2. RF time calibration
      3. Refine timewalk, etc. corrections
    3. Monitor dependence on conditions (gas mix, temperature, etc.)
    4. Calculate luminosity/normalization factor
      1. Using PS - Kei and Simon will talk to Alex about this
      2. Think up alternate method
  3. Long-term (summer)
    1. Improve FCAL/BCAL calibrations
    2. Improve detector alignment
    3. Update calibration constants and smearing in simulation
    4. Update PID calculations (requires understanding uncertainties)
    5. Calculate luminosity/normalization factor using physics channel (e.g. rho photoproduction)
      1. Requires understanding acceptances, efficiencies


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at .